Let X represent the difference between the number of heads and the number of tails obtained when a coin is tossed n times. What are the possible values of X?
Answer #1

n = number of trials

h = number of heads

t = number of tails

n = t + h

t,h = 0,1,2,.................n

X = difference of the number of heads and tails

x = h - t

x = n - t - t = n - 2t

t = 0,1,2,.............n

What are the possible values of X?

x = n - 2t

t = 0,1,2,.........n